Recycle Calculations in CHEMCAD

\underline{Problem 1}: A simplified flowsheet for the ethanol process is given in Figure 1. The specifications for the feed and each unit operation are given below:

\underline{Feed}: The feed is available at 300 k and 1 bar and has the following composition:

Ethylene 96 mol/s
Propylene 3 mol/s
Methane 1 mol/s
Water  770 mol/s

\underline{Reactor}: The reactor operates at 590 K and 69 bar. The following reactions occur:

\begin{aligned}Ethylene + W ater & → & Ethanol & conversion = 7% \\ P ropylene + W ater & → & Isopropylalcohol & conversion = 0.7% \end{aligned}

\underline{Flash}: The flash operates at a pressure of 68.5 bar and a temperature of 393 K.

\underline{Absorber}: The solvent for the absorber is pure water at a flowrate of 38 mol/s at 310 K. The vapor comes in at 393 K and the absorber operates at 68 bar.

\underline{Splitter}: The purge fraction (moles purged/moles entering the splitter) is represented by ξ.

• Determine the flowrates of each component to the reactor for a purge rate of ξ = 0.003, ξ = 0.005, and ξ = 0.007,

• Determine the mole fraction of methane at the reactor inlet for the above purge rates.

• How much ethanol goes in the absorber liquid stream if the solvent flowrate to the absorber is increased or decreased by 25% for each purge rate.

Figure 1:

All calculations should be done in CHEMCAD. Only 1 report per team needs to be submitted.